OpenStudy (anonymous):

Can you solve and explain the steps you would use in solving two step equations.

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

please do the steps like this :) and write the steps in words please! 1.... 2.... 3....

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

two step equations as in solving for a variable?

OpenStudy (anonymous):

@dumbcow @Astrophysics

OpenStudy (anonymous):

I don't know what you're getting at here. Well then, goodnight.

OpenStudy (dumbcow):

step 1: combine like terms step 2: divide both sides by "x" coefficient

OpenStudy (anonymous):

theres only two steps?

OpenStudy (anonymous):

can you give me an example? :)

OpenStudy (dumbcow):

example: 5x - 7 = 3 step 1: combine like terms ---> 5x - 7 = 3 +7 +7 ----------------------- 5x = 10 step 2: divide x = 10/5 = 2

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

how do you solve and graph inequalities? :)

OpenStudy (anonymous):

@dumbcow